A browser sells 15 grams of nuts at 0.75 pesos per grams. How many grams of nuts at one peso per gram must he sell to make his average price point at 0.85
pesos?

Answer by macston(5194)

.
x=number of grams to sell at one peso
.
(0.75 pesos/g)(15 grams)+(1 peso/g)(x grams)=(0.85 pesos/g)(x+15 grams)
(0.75)(15)+ x=(0.85x)+(0.85)(15) Subtract (0.85x+(0.75)(15)) from each side.
0.15x=(0.10)(15) Divide each side by 0.15.
x=10 ANSWER: He must sell 10 grams at 1 peso.
.
CHECK:
.
(0.75 pesos/g)15g+(1peso/g)10g=(0.85 pesos/g)25g
(11.25 pesos)+(10 pesos)=21.25 pesos
21.25 pesos=21.25 pesos